# Quench Spectroscopy for Dissipative and Non-Hermitian Quantum Lattice Models

### v1

@@ 2 Model and quench procedure
$\frac{\mathrm{d}}{\mathrm{d}t}G_{\mathbf{0}}(t)=-\sum_{\mathbf{q}\neq\mathbf{0}}\left[\left(\mathcal{G}_{\mathbf{q}}+\mathcal{H}_{\mathbf{q}}\right)F_{\mathbf{0}}(t)F_{\mathbf{q}}(t)^{*}+\mathrm{h.c}\right]-\sum_{\mathbf{q}\neq\mathbf{0}}\mathcal{F}_{\mathbf{q}}G_{\mathbf{0}}(t)G_{\mathbf{q}}(t)-2\mathcal{G}_{\mathbf{0}}G_{\mathbf{0}}(t)(G_{\mathbf{0}}(t)-1);$ (5a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{\mathbf{0}}(t)=-\mathcal{G}_{\mathbf{0}}(2G_{\mathbf{0}}(t)-3)F_{\mathbf{0}}(t)-\sum_{\mathbf{q}\neq\mathbf{0}}\mathcal{F}_{\mathbf{q}}F_{\mathbf{0}}(t)G_{\mathbf{q}}(t)-\sum_{\mathbf{q}\neq\mathbf{0}}\left(\mathcal{G}_{\mathbf{q}}+\mathcal{H}_{\mathbf{q}}\right)(2G_{\mathbf{0}}(t)+1)F_{\mathbf{q}}(t);$ (5b)
$\frac{\mathrm{d}}{\mathrm{d}t}G_{\mathbf{k}}(t)=-2\mathcal{B}_{\mathbf{k}}(t)\operatorname{Im}(F_{\mathbf{k}}(t))-\mathcal{G}_{\mathbf{k}}(F_{\mathbf{0}}(t)F_{\mathbf{k}}(t)^{*}+\mathrm{h.c})-\mathcal{F}_{\mathbf{k}}G_{\mathbf{0}}(t)G_{\mathbf{k}}(t)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathbf{k}\neq\mathbf{0};$ (5c)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{\mathbf{k}}(t)=-\left[2i\mathcal{A}_{\mathbf{k}}(t)+\mathcal{F}_{\mathbf{k}}G_{\mathbf{0}}(t)\right]F_{\mathbf{k}}(t)-\left[i\mathcal{B}_{\mathbf{k}}(t)+\mathcal{G}_{\mathbf{k}}F_{\mathbf{0}}(t)\right](2G_{\mathbf{k}}(t)+1)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathbf{k}\neq\mathbf{0},$ (5d)
where the momentum- (and possibly time-) dependent functions $\mathcal{A}_{\mathbf{k}}(t)$, $\mathcal{B}_{\mathbf{k}}(t)$, $\mathcal{F}_{\mathbf{q}}$, $\mathcal{G}_{\mathbf{q}}$ and $\mathcal{H}_{\mathbf{q}}$ are defined as follows:
$\mathcal{A}_{\mathbf{k}}(t)=4J\sum_{i=1}^{D}\sin^{2}\left(\frac{\mathbf{k}\cdot\mathbf{d}_{i}}{2}\right)+\mathcal{B}_{\mathbf{k}}(t);$ (6)
$\mathcal{B}_{\mathbf{k}}(t)=\frac{U}{L^{D}}\sum_{\mathbf{q}}G_{\mathbf{q}}(t)=U\bar{n}(t);$ (7)
$\mathcal{F}_{\mathbf{q}}=2(\mathcal{G}_{\mathbf{0}}+\mathcal{G}_{\mathbf{q}});$ (8)
$\mathcal{G}_{\mathbf{q}}=\frac{1}{L^{2D}}\sum_{\mathbf{R},\mathbf{R}^{\prime}}\gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}\cos(\mathbf{q}\cdot(\mathbf{R}-\mathbf{R}^{\prime}));$ (9)
$\mathcal{H}_{\mathbf{q}}=\frac{1}{L^{2D}}\sum_{\mathbf{R},\mathbf{R}^{\prime}}\gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}i\sin(\mathbf{q}\cdot(\mathbf{R}-\mathbf{R}^{\prime})),$ (10)
@@
$\mathcal{A}_{\mathbf{k}}(t)=4J\sin^{2}(k/2)+\mathcal{B}_{k}(t);$ (11)
$\mathcal{B}_{\mathbf{k}}(t)=\frac{U}{L}\sum_{q}G_{q}(t)=U\bar{n}(t);$ (12)
$\mathcal{F}_{\mathbf{q}}=\frac{4\gamma}{L};$ (13)
$\mathcal{G}_{\mathbf{q}}=\frac{\gamma}{L},$ (14)
$\mathcal{H}_{\mathbf{q}}=0.$ (15)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}G_{0}(t)=-\frac{\gamma}{L}\sum_{q\neq 0}(F_{0}(t)F_{q}(t)^{*}+\mathrm{h.c})-\frac{4\gamma}{L}G_{0}(t)\sum_{q\neq 0}G_{q}(t)-\frac{2\gamma}{L}G_{0}(t)\left(G_{0}(t)-1\right);$ (16a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{0}(t)=-\frac{\gamma}{L}\left(2G_{0}(t)-3\right)F_{0}(t)-\frac{4\gamma}{L}F_{0}(t)\sum_{q\neq 0}G_{q}(t)-\frac{\gamma}{L}(2G_{0}(t)+1)\sum_{q\neq 0}F_{q}(t);$ (16b)
$\frac{\mathrm{d}}{\mathrm{d}t}G_{k}(t)=-2\mathcal{B}_{k}(t)\operatorname{Im}(F_{k}(t))-\frac{\gamma}{L}\left(F_{0}(t)F_{k}(t)^{*}+\mathrm{h.c}\right)-\frac{4\gamma}{L}G_{0}(t)G_{k}(t),\leavevmode\nobreak\ \forall k\neq 0;$ (16c)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{k}(t)=-\left[2i\mathcal{A}_{k}(t)+\frac{4\gamma}{L}G_{0}(t)\right]F_{k}(t)-\left[i\mathcal{B}_{k}(t)+\frac{\gamma}{L}F_{0}(t)\right](2G_{k}(t)+1),\leavevmode\nobreak\ \forall k\neq 0,$ (16d)
@@
$G_{0}(0)=N_{0}=N-\sum_{k\neq 0}G_{k}(0);$ (17a)
$F_{0}(0)=\Theta(U)N_{0};$ (17b)
$G_{k}(0)=\frac{1}{2}\left(\frac{\mathcal{A}_{k}}{\mathcal{E}_{k}}-1\right);$ (17c)
$F_{k}(0)=-\frac{\mathcal{B}_{k}}{2\mathcal{E}_{k}},$ (17d)
@@
$\mathcal{E}_{k}=\sqrt{\mathcal{A}_{k}^{2}-\mathcal{B}_{k}^{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{A}_{k}=4J\sin^{2}(k/2)+\mathcal{B}_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}_{k}=U\bar{n}.$ (18)
3 Quench dynamics of the isolated Bose-Hubbard chain
In Ref. [59], the validity of the EoM approach to describe accurately the quench dynamics of the dissipative BH chain characterized by on-site (or long-range) two-body losses and initially confined in the SF phase has been certified against tensor networks numerical simulations. Hence, we can rely on the latter to describe the quench dynamics of the isolated BH chain in this same gapless phase for a sudden global quench on the interaction strength $U$. However, due to this new framework, the theory requires several adjustments. In what follows, we present them and benchmark the theoretical predictions of the EoM approach with those provided by a quasiparticle theory proposed in Ref. [57, 68].
In what follows, a sudden global quench on the interaction strength $U$ is performed for the isolated BH chain confined in the SF phase. To do so, the set of EoMs has to be slightly modified:
  • The quantum system being isolated has $U(1)$ symmetry, i.e. particle-number conservation, implying that the momentum- and time-dependent functions $\mathcal{A}_{k}(t)$ and $\mathcal{B}_{k}(t)$ depending on $n(t)$ remain constant in time and become equal to their initial value.
  • The coefficients $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ in the EoMs have to be replaced by $\mathcal{A}_{k,\mathrm{f}}$ and $\mathcal{B}_{k,\mathrm{f}}$ taking into account the post-quench value of $U$ denoted by $U_{\mathrm{f}}$.
  • Similarly for the initial conditions, $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ are replaced by $\mathcal{A}_{k,\mathrm{i}}$ and $\mathcal{B}_{k,\mathrm{i}}$ depending on the pre-quench value $U_{\mathrm{i}}$.
  • $\gamma=0$ for each EoM in order to remove dissipation effects.
Figure 1: Connected one-body correlation function $G_{1}(R,t)$ for a sudden global quench on $U$ of the BH chain confined in the SF phase. (top) QP approach (bottom) EoM approach. The considered parameters are: $J=1$, $U_{\mathrm{f}}=0.5$, $U_{\mathrm{i}}=1$, $N=L=400$.
@@ 3 Quench dynamics of the isolated Bose-Hubbard chain
$\frac{\mathrm{d}}{\mathrm{d}t}G_{k}(t)=-2\mathcal{B}_{k,\mathrm{f}}\operatorname{Im}(F_{k}(t));$ (19a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{k}(t)=-2i\mathcal{A}_{k,\mathrm{f}}F_{k}(t)-i\mathcal{B}_{k,\mathrm{f}}(1+2G_{k}(t)),$ (19b)
@@
$G_{k}(0)=\frac{1}{2}\left(\frac{\mathcal{A}_{k,\mathrm{i}}}{\mathcal{E}_{k,\mathrm{i}}}-1\right);$ (20)
$F_{k}(0)=-\frac{\mathcal{B}_{k,\mathrm{i}}}{2\mathcal{E}_{k,\mathrm{i}}},$ (21)
@@
$\mathcal{E}_{k,\mathrm{i}}=\sqrt{\mathcal{A}_{k,\mathrm{i}}^{2}-\mathcal{B}_{k,\mathrm{i}}^{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{A}_{k,\mathrm{i}}=4J\sin^{2}(k/2)+U_{\mathrm{i}}\bar{n},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}_{k,\mathrm{i}}=U_{\mathrm{i}}\bar{n}.$ (22)
Note that if no quench is induced, i.e. $U_{\mathrm{i}}=U_{\mathrm{f}}$, then the EoMs are strictly equal to zero and we recover $F_{k}(t)=F_{k}(0)$ and $G_{k}(t)=G_{k}(0)$ as expected. In what follows, the connected one-body correlation function $G_{1}(R,t)$ defined as $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ is considered. In reciprocal space, the latter reads:
$G_{1}(R,t)=\frac{1}{L}\sum_{k\neq 0}e^{ikR}\left(\langle\hat{n}_{k}\rangle_{t}-\langle\hat{n}_{k}\rangle_{0}\right)+\frac{1}{L}\left(\langle\hat{n}_{0}\rangle_{t}-\langle\hat{n}_{0}\rangle_{0}\right).$ (23)
According to the expression of $\hat{H}_{\mathrm{f}}$ after diagonalisation in reciprocal space, see Refs. [57, 68, 59], the latter can be simplified using the following property:
$\frac{\mathrm{d}\langle\hat{n}_{0}\rangle_{t}}{\mathrm{d}t}=i\left\langle\left[\hat{H}_{\mathrm{f}},\hat{n}_{0}\right]\right\rangle_{t}=0,$ (24)
@@
where $G_{k}(t)$ was characterised previously. Our EoM approach is compared to a quasiparticle (QP) approach based on the Heisenberg picture together with the introduction of Bogolyubov quasiparticle operators [57, 68]. Its validity has been certified numerically using tensor network based techniques  [58]. The QP approach permits to find an explicit expression of $G_{1}(R,t)$ which reads as:
$G_{1}(R,t)=\frac{1}{L}\sum_{k\neq 0}\mathcal{S}^{(1)}_{k}\cos(kR)\sin^{2}(\mathcal{E}_{k,\mathrm{f}}t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{S}^{(1)}_{k}=\frac{\left(\mathcal{A}_{k,\mathrm{i}}\mathcal{B}_{k,\mathrm{f}}-\mathcal{A}_{k,\mathrm{f}}\mathcal{B}_{k,\mathrm{i}}\right)\mathcal{B}_{k,\mathrm{f}}}{\mathcal{E}_{k,\mathrm{i}}\mathcal{E}_{k,\mathrm{f}}^{2}},$ (26)
where the amplitude function is denoted by $\mathcal{S}^{(1)}_{k}$. On Fig. 1, both theoretical approaches are compared and we find a perfect agreement between them. This validates the EoM approach defined by the set of coupled differential equations at Eq. (16). Note that both theories are equivalent since they are based on the same expression of the diagonalised form of the Hamiltonian $\hat{H}$ in Fourier space. While the EoM approach is based on the calculation of coupled first-order differential equations of quadratic correlators, the QP approach directly calculates the expression of the time-evolved Bogolyubov quasiparticle operators and thus permits to find an explicit analytical expression of the correlation functions. Regarding the density fluctuations characterised via the $G_{2}$ connected density-density correlation function, the latter may be written as follows:
Figure 2: Connected density-density correlation function $G_{2}(R,t)$ for a sudden global quench on $U$ of the BH chain confined in the SF phase. (top) QP approach (bottom) EoM approach. The considered parameters are: $J=1$, $U_{\mathrm{f}}=0.5$, $U_{\mathrm{i}}=1$, $N=L=400$.
$G_{2}(R,t)=\langle\hat{n}_{R}(t)\hat{n}_{0}(t)\rangle_{\mathrm{c}}=\langle\hat{n}_{R}\hat{n}_{0}\rangle_{t}-\langle\hat{n}_{R}\rangle_{t}\langle\hat{n}_{0}\rangle_{t}-\langle\hat{n}_{R}\hat{n}_{0}\rangle_{0}+\langle\hat{n}_{R}\rangle_{0}\langle\hat{n}_{0}\rangle_{0}.$ (27)

### v2

@@ II.1 Model and quench procedure
$\frac{\mathrm{d}}{\mathrm{d}t}G_{\mathbf{0}}(t)=$ $\leavevmode\nobreak\ -\sum_{\mathbf{q}\neq\mathbf{0}}\left[\left(\mathcal{G}_{\mathbf{q}}+\mathcal{H}_{\mathbf{q}}\right)F_{\mathbf{0}}(t)F_{\mathbf{q}}(t)^{*}+\mathrm{h.c}\right]-\sum_{\mathbf{q}\neq\mathbf{0}}\mathcal{F}_{\mathbf{q}}G_{\mathbf{0}}(t)G_{\mathbf{q}}(t)-2\mathcal{G}_{\mathbf{0}}G_{\mathbf{0}}(t)(G_{\mathbf{0}}(t)-1);$ (5a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{\mathbf{0}}(t)=$ $\leavevmode\nobreak\ -\mathcal{G}_{\mathbf{0}}(2G_{\mathbf{0}}(t)-3)F_{\mathbf{0}}(t)-\sum_{\mathbf{q}\neq\mathbf{0}}\mathcal{F}_{\mathbf{q}}F_{\mathbf{0}}(t)G_{\mathbf{q}}(t)-\sum_{\mathbf{q}\neq\mathbf{0}}\left(\mathcal{G}_{\mathbf{q}}+\mathcal{H}_{\mathbf{q}}\right)(2G_{\mathbf{0}}(t)+1)F_{\mathbf{q}}(t);$ (5b)
$\frac{\mathrm{d}}{\mathrm{d}t}G_{\mathbf{k}}(t)=$ $\leavevmode\nobreak\ -2\mathcal{B}_{\mathbf{k}}(t)\operatorname{Im}(F_{\mathbf{k}}(t))-\mathcal{G}_{\mathbf{k}}(F_{\mathbf{0}}(t)F_{\mathbf{k}}(t)^{*}+\mathrm{h.c})-\mathcal{F}_{\mathbf{k}}G_{\mathbf{0}}(t)G_{\mathbf{k}}(t)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathbf{k}\neq\mathbf{0};$ (5c)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{\mathbf{k}}(t)=$ $\leavevmode\nobreak\ -\left[2i\mathcal{A}_{\mathbf{k}}(t)+\mathcal{F}_{\mathbf{k}}G_{\mathbf{0}}(t)\right]F_{\mathbf{k}}(t)-\left[i\mathcal{B}_{\mathbf{k}}(t)+\mathcal{G}_{\mathbf{k}}F_{\mathbf{0}}(t)\right](2G_{\mathbf{k}}(t)+1)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathbf{k}\neq\mathbf{0},$ (5d)
where the momentum- and possibly time-dependent functions $\mathcal{A}_{\mathbf{k}}(t)$, $\mathcal{B}_{\mathbf{k}}(t)$, $\mathcal{F}_{\mathbf{q}}$, $\mathcal{G}_{\mathbf{q}}$ and $\mathcal{H}_{\mathbf{q}}$ are defined as follows:
$\mathcal{A}_{\mathbf{k}}(t)=4J\sum_{i=1}^{D}\sin^{2}\left(\frac{\mathbf{k}\cdot\mathbf{d}_{i}}{2}\right)+\mathcal{B}_{\mathbf{k}}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}_{\mathbf{k}}(t)=\leavevmode\nobreak\ \frac{U}{L^{D}}\sum_{\mathbf{q}}G_{\mathbf{q}}(t)=U\bar{n}(t);$ (6a)
$\mathcal{F}_{\mathbf{q}}=2(\mathcal{G}_{\mathbf{0}}+\mathcal{G}_{\mathbf{q}}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{G}_{\mathbf{q}}=\frac{1}{L^{2D}}\sum_{\mathbf{R},\mathbf{R}^{\prime}}\gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}\cos[\mathbf{q}(\mathbf{R}-\mathbf{R}^{\prime})],\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{H}_{\mathbf{q}}=\frac{i}{L^{2D}}\sum_{\mathbf{R},\mathbf{R}^{\prime}}\gamma_{||\mathbf{R}-\mathbf{R}^{\prime}||}\sin[\mathbf{q}(\mathbf{R}-\mathbf{R}^{\prime})],$ (6b)
@@
$\mathcal{A}_{\mathbf{k}}(t)=4J\sin^{2}(k/2)+\mathcal{B}_{k}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}_{\mathbf{k}}(t)=\frac{U}{L}\sum_{q}G_{q}(t)=U\bar{n}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{F}_{\mathbf{q}}=\frac{4\gamma}{L},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{G}_{\mathbf{q}}=\frac{\gamma}{L},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{H}_{\mathbf{q}}=0.$ (7a)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}G_{0}(t)=$ $\leavevmode\nobreak\ -\frac{\gamma}{L}\sum_{q\neq 0}(F_{0}(t)F_{q}(t)^{*}+\mathrm{h.c})-\frac{4\gamma}{L}G_{0}(t)\sum_{q\neq 0}G_{q}(t)-\frac{2\gamma}{L}G_{0}(t)\left(G_{0}(t)-1\right);$ (8a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{0}(t)=$ $\leavevmode\nobreak\ -\frac{\gamma}{L}\left(2G_{0}(t)-3\right)F_{0}(t)-\frac{4\gamma}{L}F_{0}(t)\sum_{q\neq 0}G_{q}(t)-\frac{\gamma}{L}(2G_{0}(t)+1)\sum_{q\neq 0}F_{q}(t);$ (8b)
$\frac{\mathrm{d}}{\mathrm{d}t}G_{k}(t)=$ $\leavevmode\nobreak\ -2\mathcal{B}_{k}(t)\operatorname{Im}(F_{k}(t))-\frac{\gamma}{L}\left(F_{0}(t)F_{k}(t)^{*}+\mathrm{h.c}\right)-\frac{4\gamma}{L}G_{0}(t)G_{k}(t),\leavevmode\nobreak\ \forall k\neq 0;$ (8c)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{k}(t)=$ $\leavevmode\nobreak\ -\left[2i\mathcal{A}_{k}(t)+\frac{4\gamma}{L}G_{0}(t)\right]F_{k}(t)-\left[i\mathcal{B}_{k}(t)+\frac{\gamma}{L}F_{0}(t)\right](2G_{k}(t)+1),\leavevmode\nobreak\ \forall k\neq 0,$ (8d)
@@
$G_{0}(0)=$ $\leavevmode\nobreak\ N_{0}=N-\sum_{k\neq 0}G_{k}(0);$ (9a)
$F_{0}(0)=$ $\leavevmode\nobreak\ \Theta(U)N_{0};$ (9b)
$G_{k}(0)=$ $\leavevmode\nobreak\ \frac{1}{2}\left(\frac{\mathcal{A}_{k}}{\mathcal{E}_{k}}-1\right);$ (9c)
$F_{k}(0)=$ $\leavevmode\nobreak\ -\frac{\mathcal{B}_{k}}{2\mathcal{E}_{k}},$ (9d)
@@
$\mathcal{E}_{k}=\sqrt{\mathcal{A}_{k}^{2}-\mathcal{B}_{k}^{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{A}_{k}=4J\sin^{2}(k/2)+\mathcal{B}_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}_{k}=U\bar{n}.$ (10)
II.2 Quench dynamics of the isolated Bose-Hubbard chain
In Ref. despresnew , the validity of the EoM approach to describe accurately the quench dynamics of the dissipative BH chain characterized by on-site (or long-range) two-body losses and initially confined in the SF-mean-field regime has been certified against tensor networks numerical simulations. Hence, we can rely on the latter to describe the quench dynamics of the isolated BH chain in this same gapless phase for a sudden global quench on the interaction strength $U$. However, due to this new framework, the theory requires several adjustments. In what follows, we present them and benchmark the theoretical predictions of the EoM approach with those provided by a quasiparticle theory proposed in Ref. despres2018 ; despres_these .
In what follows, a sudden global quench on the interaction strength $U$ is performed for the isolated BH chain confined in the SF-mean-field regime. To do so, the set of EoMs has to be slightly modified:
[…]
Figure 1: Connected one-body correlation function $G_{1}(R,t)$ for a sudden global quench on $U$ of the BH chain confined in the SF-mean-field regime. On the top panel, the QP approach is considered and on the bottom panel, the EoM approach is used. The considered dimensionless interaction parameters are: $U_{\mathrm{f}}\bar{n}/J=0.5$, $U_{\mathrm{i}}\bar{n}/J=1$.
@@ II.2 Quench dynamics of the isolated Bose-Hubbard chain
$\frac{\mathrm{d}}{\mathrm{d}t}G_{k}(t)=$ $\leavevmode\nobreak\ -2\mathcal{B}_{k,\mathrm{f}}\operatorname{Im}(F_{k}(t));$ (11a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{k}(t)=$ $\leavevmode\nobreak\ -2i\mathcal{A}_{k,\mathrm{f}}F_{k}(t)-i\mathcal{B}_{k,\mathrm{f}}(1+2G_{k}(t)),$ (11b)
@@
$G_{k}(0)=$ $\leavevmode\nobreak\ \frac{1}{2}\left(\frac{\mathcal{A}_{k,\mathrm{i}}}{\mathcal{E}_{k,\mathrm{i}}}-1\right),\leavevmode\nobreak\ \leavevmode\nobreak\ F_{k}(0)=-\frac{\mathcal{B}_{k,\mathrm{i}}}{2\mathcal{E}_{k,\mathrm{i}}},$ (12)
@@
$\mathcal{E}_{k,\mathrm{i}}=$ $\leavevmode\nobreak\ \sqrt{\mathcal{A}_{k,\mathrm{i}}^{2}-\mathcal{B}_{k,\mathrm{i}}^{2}},$ (13a)
$\mathcal{A}_{k,\mathrm{i}}=$ $\leavevmode\nobreak\ 4J\sin^{2}(k/2)+U_{\mathrm{i}}\bar{n},$ (13b)
$\mathcal{B}_{k,\mathrm{i}}=$ $\leavevmode\nobreak\ U_{\mathrm{i}}\bar{n}.$ (13c)
Note that if no quench is induced, i.e. $U_{\mathrm{i}}=U_{\mathrm{f}}$, then the EoMs are strictly equal to zero and we recover $F_{k}(t)=F_{k}(0)$ and $G_{k}(t)=G_{k}(0)$ as expected.
In what follows, the connected one-body correlation function $G_{1}(R,t)$ defined as $G_{1}(R,t)=\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{t}-\langle\hat{b}^{{\dagger}}_{R}\hat{b}_{0}\rangle_{0}$ is considered. In reciprocal space, the latter reads:
$G_{1}(R,t)=\frac{1}{L}\sum_{k}e^{ikR}\left(\langle\hat{n}_{k}\rangle_{t}-\langle\hat{n}_{k}\rangle_{0}\right).$ (14)
The latter can be further simplified. The expression of $\hat{H}_{\mathrm{f}}$ in reciprocal space is given at Eq. (3c) where $\mathcal{A}_{k}(t)$ and $\mathcal{B}_{k}(t)$ are respectively replaced by $\mathcal{A}_{k,\mathrm{f}}$ and $\mathcal{B}_{k,\mathrm{f}}$. The time dependence vanishes due to the particle-number conservation since the isolated BH chain is considered, see also Refs. despres2018 ; despres_these ; despresnew . From the latter, we notice that the condensate mode $k=0$ has been integrated out using the standard mean-field approximation and thus leading to the filling (or equivalently the density) dependence $\bar{n}$ within the post-quench momentum-dependent functions $\mathcal{A}_{k,\mathrm{f}}$ and $\mathcal{B}_{k,\mathrm{f}}$. It immediately follows the property:
$\frac{\mathrm{d}\langle\hat{n}_{0}\rangle_{t}}{\mathrm{d}t}=i\langle[\hat{H}_{\mathrm{f}},\hat{n}_{0}]\rangle_{t}=0,$ (15)
@@
where $G_{k}(t)=\langle\hat{n}_{k}\rangle_{t}=\langle\hat{b}^{{\dagger}}_{k}\hat{b}_{k}\rangle_{t}$. Our EoM approach is compared to a quasiparticle (QP) approach based on the Heisenberg picture together with the introduction of Bogolyubov quasiparticle operators despres2018 ; despres_these . Its validity has been certified numerically using tensor network based techniques  despres2019 . The QP approach permits to find an explicit expression of $G_{1}(R,t)$ which reads as:
$G_{1}(R,t)=$ $\leavevmode\nobreak\ \frac{1}{L}\sum_{k\neq 0}\mathcal{S}^{(1)}_{k}\cos(kR)\sin^{2}(\mathcal{E}_{k,\mathrm{f}}t),$ (17)
where the amplitude function is denoted by $\mathcal{S}^{(1)}_{k}$ and defined as:
$\mathcal{S}^{(1)}_{k}=$ $\leavevmode\nobreak\ \frac{\left(\mathcal{A}_{k,\mathrm{i}}\mathcal{B}_{k,\mathrm{f}}-\mathcal{A}_{k,\mathrm{f}}\mathcal{B}_{k,\mathrm{i}}\right)\mathcal{B}_{k,\mathrm{f}}}{\mathcal{E}_{k,\mathrm{i}}\mathcal{E}_{k,\mathrm{f}}^{2}}.$ (18)
On Fig. 1, both theoretical approaches are compared and we find a perfect agreement between them. This validates the EoM approach defined by the set of coupled differential equations at Eq. (8). Note that both theories are equivalent since they are based on the same expression of the diagonalized form of the Hamiltonian $\hat{H}$ in Fourier space. While the EoM approach is based on the calculation of coupled first-order differential equations of quadratic correlators, the QP approach directly calculates the expression of the time-evolved Bogolyubov quasiparticle operators and thus permits to find an explicit analytical expression of the correlation functions.
Regarding the density fluctuations characterized via the $G_{2}$ connected density-density correlation function, the latter may be written as follows:
Figure 2: Connected density-density correlation function $G_{2}(R,t)$ for a sudden global quench on $U$ of the BH chain confined in the SF-mean-field regime. On the top panel, the QP approach is considered and on the bottom panel, the EoM approach is used. The considered dimensionless interaction parameters are: $U_{\mathrm{f}}\bar{n}/J=0.5$, $U_{\mathrm{i}}\bar{n}/J=1$.
$G_{2}(R,t)=\langle\hat{n}_{R}\hat{n}_{0}\rangle_{t}-\langle\hat{n}_{R}\rangle_{t}\langle\hat{n}_{0}\rangle_{t}-\langle\hat{n}_{R}\hat{n}_{0}\rangle_{0}+\langle\hat{n}_{R}\rangle_{0}\langle\hat{n}_{0}\rangle_{0}.$ (19)